OpenStudy (anonymous):

The numerator of a fraction is 5 times the denominator. If 3 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 17/4. Find the original fraction

OpenStudy (darkprince14):

Let x represent the denominator... Then do your work from there...

OpenStudy (anonymous):

idont know how thats why im asking

OpenStudy (darkprince14):

oh.. so the original fraction would be 5x/ x since the numerator is 5 times the denominator... If 3 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 17/4. \[\frac{ 5x-3 }{ x+3 } = \frac{ 17 }{ 4 }\] Cross Multiplication \[4(5x -3) = 17(x+3)\] Simplify.. Can you do this from here and find x??

OpenStudy (darkprince14):

@fenton24

OpenStudy (anonymous):

well x= 21 but im looking for the original fraction, so it still doenst solve it

OpenStudy (darkprince14):

If you look at the original fraction with the x part.. It looks like this\[\frac{ 5x }{ x }\]in fractions, if you have a similar factor in the numerator and the denominator, you may cancel it because it is equal to 1.\[\frac{ 5x }{ x } = 5\]Then the fraction will become 5